# Estimating the FDR of variable selection

### v2

@@ Appendix D Deferred proofs
$\mathcal{R}_{k}$ depends on $\boldsymbol{Y}$ only through $\boldsymbol{X}_{B_{k}}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y}$. The p-value $p_{j}$ depends on $\boldsymbol{Y}$ only through $(\boldsymbol{X}_{B_{k}}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y},\textnormal{RSS})$, where $\textnormal{RSS}=\left\|\boldsymbol{Y}\right\|^{2}-\left\|\boldsymbol{X}(\boldsymbol{X}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{X})^{-1}\boldsymbol{X}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y}\right\|^{2}$.
@@
Let $\boldsymbol{Y}^{(k)}$ be the projection of $\boldsymbol{Y}$ onto the subspace spanned by the columns of $\boldsymbol{X}_{B_{k}}$, and denote the projection of $\boldsymbol{Y}$ onto the columns of $\boldsymbol{X}$ as $\widehat{\boldsymbol{Y}}=\boldsymbol{X}(\boldsymbol{X}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{X})^{-1}\boldsymbol{X}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y}$. Then
$\widehat{\boldsymbol{Y}}=\sum_{k=1}^{K}\boldsymbol{Y}^{(k)}$
and the $\boldsymbol{Y}^{(k)}$s are orthogonal to each other under Assumption 3.1. We rewrite the Lasso coefficients
@@
$=\operatorname*{argmin}_{\boldsymbol{t}}\;\frac{1}{2}\,\left\|\sum_{k=1}^{K}\boldsymbol{Y}^{(k)}-\boldsymbol{X}_{B_{k}}\boldsymbol{t}_{B_{k}}\right\|_{2}^{2}+\lambda\sum_{k=1}^{K}\left\|\boldsymbol{t}_{B_{k}}\right\|_{1}$
$=\operatorname*{argmin}_{\boldsymbol{t}}\;\sum_{k=1}^{K}\frac{1}{2}\,\left\|\boldsymbol{Y}^{(k)}-\boldsymbol{X}_{B_{k}}\boldsymbol{t}_{B_{k}}\right\|_{2}^{2}+\lambda\left\|\boldsymbol{t}_{B_{k}}\right\|_{1}.$
We see solving for $\widehat{\theta}^{\textnormal{Lasso}}$ is equivalent to solving for Lasso coefficients for $K$ subproblems of regressing $\boldsymbol{Y}$ on $\boldsymbol{X}_{B_{k}}$. So $\widehat{\theta}^{\textnormal{Lasso}}_{B_{k}}(\lambda)$ depends on $\boldsymbol{Y}$ only through $\boldsymbol{X}_{B_{k}}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y}^{(k)}$, which is equal to $\boldsymbol{X}_{B_{k}}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y}$. Therefore, $\mathcal{R}_{k}=\left\{j\in B_{k}:\;\widehat{\theta}^{\textnormal{Lasso}}_{j}(\lambda)\neq 0\right\}$ depends on $\boldsymbol{Y}$ only through $\boldsymbol{X}_{B_{k}}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y}$.
@@
$\left((\boldsymbol{X}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{X})^{-1}\right)_{j}\cdot\boldsymbol{X}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y}=\left((\boldsymbol{X}_{B_{k}}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{X}_{B_{k}})^{-1}\right)_{i(j,k)}\cdot\boldsymbol{X}_{B_{k}}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y},$
where $i(j,k)$ is the index of variable $X_{j}$ in block $B_{k}$. Therefore $p_{j}$ depends on $\boldsymbol{Y}$ only through $(\boldsymbol{X}_{B_{k}}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y},\textnormal{RSS})$. ∎
@@
Recall in the Gaussian linear model, we conditional on the explanatory variables $X$ and treat them as fixed. By Lemma D.1 and the block-orthogonality, $R_{k}$ are independent for $k=1,\ldots,K$. Our goal is to show $R=\sum_{k}R_{k}$ is stochastically larger than the sum of certain independent Bernoulli random variables. Then the proposition follows by the concentration laws.
Part 1. By Lemma D.1 and the first order condition, $R_{k}\geq 1$ if and only if there exist some $j\in B_{k}$ such that $|\boldsymbol{X}_{j}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y}|\geq\lambda$. We have
$\mathbb{P}(R_{k}\geq 1)=\mathbb{P}(\bigcup_{j\in B_{k}}\left\{|\boldsymbol{X}_{j}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y}|\geq\lambda\right\})\geq\mathbb{P}(|\boldsymbol{X}_{j}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y}|\geq\lambda)\geq 2(1-\Phi(\lambda/\sigma)).$
@@
As a consequence, $R=\sum_{k=1}^{K}R_{k}$ is stochastically larger than $\sum_{k=1}^{K}A_{k}$, where $A_{k}$ are $K$ independent Bernoulli random variables with success probability $\delta$. By the multiplicative Chernoff bound, we have
$\mathbb{P}(R\leq Kq)\;\leq\;\mathbb{P}\left(\sum_{k=1}^{K}A_{k}\leq\left(1-\frac{\delta-q}{\delta}\right)K\delta\right)\;\leq\;\exp\left(-\frac{(\delta-q)^{2}}{2\delta}\cdot K\right).$
Part 2. To avoid triviality, assume $K\xi$ is an integer. Let $k$ denote any one of the $K\xi$ blocks that contains a variable with $|\boldsymbol{X}_{j}\cdot\boldsymbol{\mu}|\geq\lambda$. Note $\boldsymbol{X}_{j}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y}$ is Gaussian with mean $\boldsymbol{X}_{j}\cdot\boldsymbol{\mu}$. By a similar reasoning as in Part 1,
$\mathbb{P}(R_{k}\geq 1)\;\geq\;\mathbb{P}(|\boldsymbol{X}_{j}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y}|\geq\lambda)\;\geq\;\frac{1}{2}.$
As a consequence, $R=\sum_{k=1}^{K}R_{k}$ is stochastically larger than $\sum_{k=1}^{K\xi}A_{k}$, where $A_{k}$ are $K\xi$ independent Bernoulli random variables with success probability $1/2$. By the multiplicative Chernoff bound, we have
$\mathbb{P}(R\leq Kq)\;\leq\;\mathbb{P}\left(\sum_{k=1}^{K\xi}A_{k}\leq\left(1-\frac{\xi-2q}{\xi}\right)\frac{K\xi}{2}\right)\;\leq\;\exp\left(-\frac{(\xi-2q)^{2}}{4\xi}\cdot K\right).$
@@
Notice in (2), if we scale both $\boldsymbol{Y}$ and $\lambda$ by $1/\sigma$, the optimal solution $\boldsymbol{\widehat{\theta}^{\textnormal{Lasso}}}$ will be scaled by $1/\sigma$ as well, but the selected variables do not change. Hence in the proof, we assume $\sigma^{2}=1$ without loss of generality.
Let $\boldsymbol{X}_{B_{k}}$ be the columns of $\boldsymbol{X}$ whose indices are in $B_{k}$ and $\boldsymbol{X}_{-B_{k}}$ be its complement. Let $\boldsymbol{X}_{B_{k}-j}$ be the same as $\boldsymbol{X}_{B_{k}}$ but with $\boldsymbol{X}_{j}$ excluded.
@@
$\widehat{\textnormal{FDR}}^{*}_{j}$ $=\mathbb{E}_{H_{j}}\left[\frac{\mathbf{1}\left\{j\in\mathcal{R}\right\}}{|\mathcal{R}\cup\left\{j\right\}|}\mid\boldsymbol{S}_{j}\right]=\mathbb{E}_{H_{j}}\left[\frac{\mathbf{1}\left\{j\in\mathcal{R}_{k}\right\}}{|\mathcal{R}_{k}\cup\left\{j\right\}|+\sum_{l\neq k}R_{l}}\mid\boldsymbol{X}_{B_{k}-j}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y},\boldsymbol{X}_{-B_{k}}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y},r_{j}^{2}\right],$
where $r_{j}^{2}=\textnormal{RSS}+u_{j}^{2}$ and $u_{j}$ is the norm of the projection of $\boldsymbol{Y}$ on to the direction of $(\boldsymbol{I}-\boldsymbol{X}_{-j}(\boldsymbol{X}_{-j}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{X}_{-j})^{-1}\boldsymbol{X}_{-j}^{\mkern-1.5mu\mathsf{T}})\boldsymbol{X}_{j}$. By Lemma D.1, $\sum_{l\neq k}R_{l}$ depends on $\boldsymbol{Y}$ only through $\boldsymbol{X}_{-B_{k}}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y}$ and $\mathcal{R}_{k}$ depends on $\boldsymbol{Y}$ only through $\boldsymbol{X}_{B_{k}}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y}$. The conditional expectation $\widehat{\textnormal{FDR}}^{*}_{j}$ is an one-dimensional integration over $u_{j}$. To avoid confusion, we use $u_{j}$ to denote the value corresponding to the observed $\boldsymbol{Y}$ and $v_{j}$ to denote a placeholder for $u_{j}$ in the integral. We write
$\widehat{\textnormal{FDR}}^{*}_{j}=\int_{\left\{v_{j}:\;j\in\mathcal{R}_{k}\right\}}\frac{1}{f(v_{j};\boldsymbol{X}_{B_{k}-j}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y},\boldsymbol{X}_{-B_{k}}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y})}\,\mu_{j}(dv_{j}\mid\boldsymbol{S}_{j}),$
@@
$f(v_{j};\boldsymbol{X}_{B_{k}-j}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y},\boldsymbol{X}_{-B_{k}}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y})=|\mathcal{R}_{k}\cup\left\{j\right\}|+\sum_{l\neq k}R_{l}$
@@
$\mu_{j}(A\mid\boldsymbol{S}_{j})=\mathbb{P}_{H_{j}}(v_{j}\in A\mid\boldsymbol{X}_{B_{k}-j}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y},\boldsymbol{X}_{-B_{k}}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y},r_{j}^{2})=\mathbb{P}_{H_{j}}(v_{j}\in A\mid\boldsymbol{X}_{B_{k}-j}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y},r_{j}^{2}).$
We suppress $\boldsymbol{S}_{j}$ and write $\mu_{j}(A)$ for simplicity of notation. Note the measure $\mu_{j}$ depends on $\boldsymbol{Y}$ only through $(\boldsymbol{X}_{B_{k}-j}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y},r_{j}^{2})$. Note
$f(v_{j})\in\left[1+\sum_{l\neq k}R_{l},\;b+\sum_{l\neq k}R_{l}\right]\subseteq\left[1\vee\left(-b+\sum_{k=1}^{K}R_{k}\right),\;b+\sum_{k=1}^{K}R_{k}\right].$
By Lemma D.2, for some $\gamma_{j}\in[-b,b]$, we have
$\widehat{\textnormal{FDR}}^{*}_{j}=\frac{\mu_{j}(j\in\mathcal{R}_{k})}{1\vee(R+\gamma_{j})}.$
@@
$\left|\widehat{\textnormal{FDR}}^{*}_{j}-\frac{\mu_{j}(j\in\mathcal{R}_{k})}{1\vee R}\right|=\left|\frac{\mu_{j}(j\in\mathcal{R}_{k})}{1\vee(R+\gamma_{j})}-\frac{\mu_{j}(j\in\mathcal{R}_{k})}{1\vee R}\right|\leq\frac{b}{1\vee R^{2}}\cdot\frac{1\vee R}{1\vee(R-b)}\leq\frac{4b^{2}}{1\vee R^{2}}.$
For notational convenience, we suppress taking maximum with $1$ for $R$. Through out the section, we have $R$ represents $1\vee R$ and $\sum_{k}R_{k}$ represents $1\vee\sum_{k}R_{k}$. Then we can write
$\widehat{\textnormal{FDR}}^{*}_{j}=\frac{\mu_{j}(j\in\mathcal{R}_{k})}{\sum_{k}R_{k}}+O\left(\frac{b^{2}}{R^{2}}\right).$
@@
$T_{k}=\sum_{j\in B_{k}}\mu_{j}(j\in\mathcal{R}_{k})\cdot\frac{\psi_{j}(p_{j})}{\mathbb{E}[\psi_{j}(p_{j})\mid\boldsymbol{S}_{j}]}.$
@@
$\sum_{j\in B_{k}}\widehat{\textnormal{FDR}}_{j}=\frac{T_{k}}{\sum_{k}R_{k}}+O\left(\frac{c_{\phi}b^{3}}{R^{2}}\right)$
@@
$\widehat{\textnormal{FDR}}=\sum_{k=1}^{K}\sum_{j\in B_{k}}\widehat{\textnormal{FDR}}_{j}=\frac{\sum_{k}T_{k}}{\sum_{k}R_{k}}+O\left(\frac{Kc_{\phi}b^{3}}{R^{2}}\right)$
Note we can compute $(\boldsymbol{X}_{B_{k}-j}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y},r_{j}^{2})$ from $(\boldsymbol{X}_{B_{k}}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y},\textnormal{RSS})$ for any $j\in B_{k}$. Moreover, by Lemma D.1, for any $j\in B_{k}$, $p_{j}$ depends on $Y$ only through $(\boldsymbol{X}_{B_{k}}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y},\textnormal{RSS})$. So $T_{k}$ depends on $Y$ only through $(\boldsymbol{X}_{B_{k}}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y},\textnormal{RSS})$. Therefore, condition on RSS, both $\left\{T_{k}:\;k\in[K]\right\}$ and $\left\{R_{k}:\;k\in[K]\right\}$ are sets of conditional independent random variables respectively. We denote $T=\sum_{k}T_{k}$ for simplicity.
@@
Step 1. $\textnormal{Var}\left({\sum_{k}T_{k}}/{\sum_{k}R_{k}}\mid\textnormal{RSS}\right)\leq C(c_{\phi}^{2}b^{4}/Kq^{4}+c_{\phi}^{2}b^{2}K^{2}\rho)$ almost surely for some $C>0$ that does not depend on RSS.
[…]
@@
$=$ $\frac{1}{2}\,\mathbb{E}_{\textnormal{RSS}}\left[\left(\frac{T}{R}-\frac{T^{\prime}}{R^{\prime}}\right)^{2}\mid R,R^{\prime}>Kq\right]\cdot\mathbb{P}_{\textnormal{RSS}}(R,R^{\prime}>Kq)$
$+\frac{1}{2}\,\mathbb{E}_{\textnormal{RSS}}\left[\left(\frac{T}{R}-\frac{T^{\prime}}{R^{\prime}}\right)^{2}\mid R\wedge R^{\prime}\leq Kq\right]\cdot\mathbb{P}_{\textnormal{RSS}}(R\wedge R^{\prime}\leq Kq),$

### v4

@@ Appendix D Variance of FDR estimation
$\mathcal{R}_{b}$ depends on $\boldsymbol{Y}$ only through $\boldsymbol{X}_{(b)}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y}$. The p-value $p_{j}$ depends on $\boldsymbol{Y}$ only through $(\boldsymbol{X}_{(b)}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y},\textnormal{RSS})$, where $\textnormal{RSS}=\left\|\boldsymbol{Y}\right\|^{2}-\left\|\boldsymbol{X}(\boldsymbol{X}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{X})^{-1}\boldsymbol{X}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y}\right\|^{2}$.
@@
Let $\boldsymbol{Y}^{(b)}$ be the projection of $\boldsymbol{Y}$ onto the subspace spanned by the columns of $\boldsymbol{X}_{(b)}$, and denote the projection of $\boldsymbol{Y}$ onto the columns of $\boldsymbol{X}$ as $\widehat{\boldsymbol{Y}}=\boldsymbol{X}(\boldsymbol{X}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{X})^{-1}\boldsymbol{X}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y}$. Then
$\widehat{\boldsymbol{Y}}=\sum_{b=1}^{B}\boldsymbol{Y}^{(b)}$
and the $\boldsymbol{Y}^{(b)}$s are orthogonal to each other under Assumption 3.1. We rewrite the Lasso coefficients
@@
$=\operatorname*{argmin}_{\boldsymbol{t}}\;\frac{1}{2}\,\left\|\sum_{b=1}^{B}\boldsymbol{Y}^{(b)}-\boldsymbol{X}_{(b)}\boldsymbol{t}_{(b)}\right\|_{2}^{2}+\lambda\sum_{b=1}^{B}\left\|\boldsymbol{t}_{(b)}\right\|_{1}$
$=\operatorname*{argmin}_{\boldsymbol{t}}\;\sum_{b=1}^{B}\frac{1}{2}\,\left\|\boldsymbol{Y}^{(b)}-\boldsymbol{X}_{(b)}\boldsymbol{t}_{(b)}\right\|_{2}^{2}+\lambda\left\|\boldsymbol{t}_{(b)}\right\|_{1}.$
We see solving for $\widehat{\theta}^{\textnormal{Lasso}}$ is equivalent to solving for Lasso coefficients for $B$ subproblems of regressing $\boldsymbol{Y}$ on $\boldsymbol{X}_{(b)}$. So $\widehat{\theta}^{\textnormal{Lasso}}_{(b)}(\lambda)$ depends on $\boldsymbol{Y}$ only through $\boldsymbol{X}_{(b)}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y}^{(b)}$, which is equal to $\boldsymbol{X}_{(b)}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y}$. Therefore, $\mathcal{R}_{b}=\left\{j\in(b):\;\widehat{\theta}^{\textnormal{Lasso}}_{j}(\lambda)\neq 0\right\}$ depends on $\boldsymbol{Y}$ only through $\boldsymbol{X}_{(b)}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y}$.
@@
$\left((\boldsymbol{X}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{X})^{-1}\right)_{j}\cdot\boldsymbol{X}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y}=\left((\boldsymbol{X}_{(b)}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{X}_{(b)})^{-1}\right)_{i(j,b)}\cdot\boldsymbol{X}_{(b)}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y},$
where $i(j,b)$ is the index of variable $X_{j}$ in block $(b)$. Therefore $p_{j}$ depends on $\boldsymbol{Y}$ only through $(\boldsymbol{X}_{(b)}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y},\textnormal{RSS})$. ∎
@@
Recall in the Gaussian linear model, we condition on the explanatory variables $\boldsymbol{X}$ and treat them as fixed. By Lemma D.1 and the block-orthogonality, $R_{b}$ are independent for $b=1,\ldots,B$. Our goal is to show $R=\sum_{b}R_{b}$ is stochastically larger than the sum of certain independent Bernoulli random variables. Then the proposition follows by the concentration laws.
Part 1. By Lemma D.1 and the first order condition, $R_{b}\geq 1$ if and only if there exist some $j\in(b)$ such that $|\boldsymbol{X}_{j}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y}|\geq\lambda$. We have
$\mathbb{P}(R_{b}\geq 1)=\mathbb{P}(\bigcup_{j\in(b)}\left\{|\boldsymbol{X}_{j}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y}|\geq\lambda\right\})\geq\mathbb{P}(|\boldsymbol{X}_{j}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y}|\geq\lambda)\geq 2(1-\Phi(\lambda/\sigma)).$
@@
As a consequence, $R=\sum_{b=1}^{B}R_{b}$ is stochastically larger than $\sum_{b=1}^{B}A_{b}$, where $A_{b}$ are $B$ independent Bernoulli random variables with success probability $\delta$. By the multiplicative Chernoff bound, we have
$\mathbb{P}(R\leq Bq)\;\leq\;\mathbb{P}\left(\sum_{b=1}^{B}A_{b}\leq\left(1-\frac{\delta-q}{\delta}\right)B\delta\right)\;\leq\;\exp\left(-\frac{(\delta-q)^{2}}{2\delta}\cdot B\right).$
Part 2. To avoid triviality, assume $B\xi$ is an integer. Let $b$ denote any one of the $B\xi$ blocks that contains a variable with $|\boldsymbol{X}_{j}\cdot\boldsymbol{\mu}|\geq\lambda$. Note $\boldsymbol{X}_{j}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y}$ is Gaussian with mean $\boldsymbol{X}_{j}\cdot\boldsymbol{\mu}$. By a similar reasoning as in Part 1,
$\mathbb{P}(R_{b}\geq 1)\;\geq\;\mathbb{P}(|\boldsymbol{X}_{j}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y}|\geq\lambda)\;\geq\;\frac{1}{2}.$
As a consequence, $R=\sum_{b=1}^{B}R_{b}$ is stochastically larger than $\sum_{b=1}^{B\xi}A_{b}$, where $A_{b}$ are $B\xi$ independent Bernoulli random variables with success probability $1/2$. By the multiplicative Chernoff bound, we have
$\mathbb{P}(R\leq Bq)\;\leq\;\mathbb{P}\left(\sum_{b=1}^{B\xi}A_{b}\leq\left(1-\frac{\xi-2q}{\xi}\right)\frac{B\xi}{2}\right)\;\leq\;\exp\left(-\frac{(\xi-2q)^{2}}{4\xi}\cdot B\right).$
@@
Notice in (2), if we scale both $\boldsymbol{Y}$ and $\lambda$ by $1/\sigma$, the optimal solution $\boldsymbol{\widehat{\theta}^{\textnormal{Lasso}}}$ will be scaled by $1/\sigma$ as well, but the selected variables do not change. Hence in the proof, we assume $\sigma^{2}=1$ without loss of generality. Dependence on $\sigma^{2}$ can be restored by rescaling $\boldsymbol{Y}$ and $\lambda$.
Let $\boldsymbol{X}_{(b)}$ be the columns of $\boldsymbol{X}$ whose indices are in $(b)$ and $\boldsymbol{X}_{-(b)}$ be its complement. Let $\boldsymbol{X}_{(b);-j}$ be the same matrix as $\boldsymbol{X}_{(b)}$ but with $\boldsymbol{X}_{j}$ excluded.
@@
$\widehat{\textnormal{FDR}}^{*}_{j}$ $=\mathbb{E}_{H_{j}}\left[\frac{\mathbf{1}\left\{j\in\mathcal{R}\right\}}{|\mathcal{R}\cup\left\{j\right\}|}\mid\boldsymbol{S}_{j}\right]=\mathbb{E}_{H_{j}}\left[\frac{\mathbf{1}\left\{j\in\mathcal{R}_{b}\right\}}{|\mathcal{R}_{b}\cup\left\{j\right\}|+\sum_{l\neq b}R_{l}}\mid\boldsymbol{X}_{(b);-j}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y},\boldsymbol{X}_{-(b)}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y},r_{j}^{2}\right],$
where $r_{j}^{2}=\textnormal{RSS}+u_{j}^{2}$ and $u_{j}$ is the norm of the projection of $\boldsymbol{Y}$ on to the direction of $(\boldsymbol{I}-\boldsymbol{X}_{-j}(\boldsymbol{X}_{-j}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{X}_{-j})^{-1}\boldsymbol{X}_{-j}^{\mkern-1.5mu\mathsf{T}})\boldsymbol{X}_{j}$. By Lemma D.1, $\sum_{l\neq b}R_{l}$ depends on $\boldsymbol{Y}$ only through $\boldsymbol{X}_{-(b)}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y}$ and $\mathcal{R}_{b}$ depends on $\boldsymbol{Y}$ only through $\boldsymbol{X}_{(b)}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y}$. The conditional expectation $\widehat{\textnormal{FDR}}^{*}_{j}$ is a one-dimensional integration over $u_{j}$. To avoid confusion, we use $u_{j}$ to denote the value corresponding to the observed $\boldsymbol{Y}$ and $v_{j}$ to denote a placeholder for $u_{j}$ in the integral. We write
$\widehat{\textnormal{FDR}}^{*}_{j}=\int_{\left\{v_{j}:\;j\in\mathcal{R}_{b}\right\}}\frac{1}{f(v_{j};\boldsymbol{X}_{(b);-j}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y},\boldsymbol{X}_{-(b)}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y})}\,\mu_{j}(dv_{j}\mid\boldsymbol{S}_{j}),$
@@
$f(v_{j};\boldsymbol{X}_{(b);-j}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y},\boldsymbol{X}_{-(b)}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y})=|\mathcal{R}_{b}\cup\left\{j\right\}|+\sum_{l\neq b}R_{l}$
@@
$\mu_{j}(A\mid\boldsymbol{S}_{j})=\mathbb{P}_{H_{j}}(v_{j}\in A\mid\boldsymbol{X}_{(b);-j}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y},\boldsymbol{X}_{-(b)}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y},r_{j}^{2})=\mathbb{P}_{H_{j}}(v_{j}\in A\mid\boldsymbol{X}_{(b);-j}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y},r_{j}^{2}).$
We suppress $\boldsymbol{S}_{j}$ and write $\mu_{j}(A)$ for simplicity of notation. Note the measure $\mu_{j}$ depends on $\boldsymbol{Y}$ only through $(\boldsymbol{X}_{(b);-j}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y},r_{j}^{2})$. Note
$f(v_{j})\in\left[1+\sum_{l\neq b}R_{l},\;m+\sum_{l\neq b}R_{l}\right]\subseteq\left[1\vee\left(-m+\sum_{b=1}^{B}R_{b}\right),\;m+\sum_{b=1}^{B}R_{b}\right].$
By Lemma D.2, for some $\gamma_{j}\in[-m,m]$, we have
$\widehat{\textnormal{FDR}}^{*}_{j}=\frac{\mu_{j}(j\in\mathcal{R}_{b})}{1\vee(R+\gamma_{j})}.$
@@
$\left|\widehat{\textnormal{FDR}}^{*}_{j}-\frac{\mu_{j}(j\in\mathcal{R}_{b})}{1\vee R}\right|=\left|\frac{\mu_{j}(j\in\mathcal{R}_{b})}{1\vee(R+\gamma_{j})}-\frac{\mu_{j}(j\in\mathcal{R}_{b})}{1\vee R}\right|\leq\frac{m}{1\vee R^{2}}\cdot\frac{1\vee R}{1\vee(R-m)}\leq\frac{4m^{2}}{1\vee R^{2}}.$
For notational convenience, we suppress taking maximum with $1$ for $R$. Throughout the section, we have $R$ represents $1\vee R$ and $\sum_{b}R_{b}$ represents $1\vee\sum_{b}R_{b}$. Then we can write
$\widehat{\textnormal{FDR}}^{*}_{j}=\frac{\mu_{j}(j\in\mathcal{R}_{b})}{\sum_{b}R_{b}}+O\left(\frac{m^{2}}{R^{2}}\right).$
@@
$T_{b}=\sum_{j\in(b)}\mu_{j}(j\in\mathcal{R}_{b})\cdot\frac{\psi_{j}(p_{j})}{\mathbb{E}[\psi_{j}(p_{j})\mid\boldsymbol{S}_{j}]}.$
@@
$\sum_{j\in(b)}\widehat{\textnormal{FDR}}_{j}=\frac{T_{b}}{\sum_{b}R_{b}}+O\left(\frac{c_{\phi}m^{3}}{R^{2}}\right)$
@@
$\widehat{\textnormal{FDR}}=\sum_{b=1}^{B}\sum_{j\in(b)}\widehat{\textnormal{FDR}}_{j}=\frac{\sum_{b}T_{b}}{\sum_{b}R_{b}}+O\left(\frac{Bc_{\phi}m^{3}}{R^{2}}\right)$
Note we can compute $(\boldsymbol{X}_{(b);-j}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y},r_{j}^{2})$ from $(\boldsymbol{X}_{(b)}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y},\textnormal{RSS})$ for any $j\in(b)$. Moreover, by Lemma D.1, for any $j\in(b)$, $p_{j}$ depends on $Y$ only through $(\boldsymbol{X}_{(b)}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y},\textnormal{RSS})$. So $T_{b}$ depends on $Y$ only through $(\boldsymbol{X}_{(b)}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y},\textnormal{RSS})$. Therefore, condition on RSS, both $\left\{T_{b}:\;b\in[B]\right\}$ and $\left\{R_{b}:\;b\in[B]\right\}$ are sets of conditional independent random variables respectively. We denote $T=\sum_{b}T_{b}$ for simplicity.
@@
Step 1. $\textnormal{Var}\left({\sum_{b}T_{b}}/{\sum_{b}R_{b}}\mid\textnormal{RSS}\right)\leq C(c_{\phi}^{2}m^{4}/Bq^{4}+c_{\phi}^{2}m^{2}B^{2}\rho)$ almost surely for some $C>0$ that does not depend on RSS.
For simplicity of notation, we denote $\textnormal{Var}(\cdot|\cdot,\textnormal{RSS})$ as $\textnormal{Var}_{\textnormal{RSS}}(\cdot|\cdot)$ and $\mathbb{E}(\cdot|\cdot,\textnormal{RSS})$ as $\mathbb{E}_{\textnormal{RSS}}(\cdot|\cdot)$. Let $q$ be the number in Proposition 3.1. We have
@@
$=$ $\frac{1}{2}\,\mathbb{E}_{\textnormal{RSS}}\left[\left(\frac{T}{R}-\frac{T^{\prime}}{R^{\prime}}\right)^{2}\mid R,R^{\prime}>Bq\right]\cdot\mathbb{P}_{\textnormal{RSS}}(R,R^{\prime}>Bq)$
$+\frac{1}{2}\,\mathbb{E}_{\textnormal{RSS}}\left[\left(\frac{T}{R}-\frac{T^{\prime}}{R^{\prime}}\right)^{2}\mid R\wedge R^{\prime}\leq Bq\right]\cdot\mathbb{P}_{\textnormal{RSS}}(R\wedge R^{\prime}\leq Bq),$